OpenStudy (whynot):

During a certain 21 hour period, the temperature at time t was T(t)=49+6t-1/3t^2. What was the average temperature during that period? During a certain 21 hour period, the temperature at time t was T(t)=49+6t-1/3t^2. What was the average temperature during that period? @Mathematics

OpenStudy (anonymous):

T(0) = 49 T(21) = 49+6(21) -1/3(21)^2 ave = (T(21)-T(0))/21

OpenStudy (anonymous):

you can work this out

OpenStudy (anonymous):

actually should be t(0) - t(21)/(0-21) = (49-28)/(0-21) =-1 degrees/hour

OpenStudy (anonymous):

i was too quick on the first answer i gave

OpenStudy (whynot):

Thank you